What is the total charge of all the electrons in 9.11 L of liquid water?
Answer #1
  1. The molar mass of an H2OH2O molecule is 2×1+1×16=182×1+1×16=18 grams
  2. The total number of electrons per H2OH2O molecule is 2×1+1×8=102×1+1×8=10
  3. The number of moles of H2OH2O in 1 kg of water is 1000/18=55.555...1000/18=55.555...
  4. The number of moles of H2OH2O in 9.1 kg of water is =9.1*55.5555=505.555
  5. The number of molecules of H2OH2O in 1 kg of water is 505.555...×6.022×1023=31.84996815...×1025
  6. Therefore, the number of electrons in 9.1 kg of water is31.84996815×1025×10*9.1=289.744×1026
  7. Finally, the total charge of those electrons is 289.744×1026×−1.602176×10−19= -483.9304 ×107Coulombs.